MONT 101N – Analyzing Environmental Data Paper Assignment – Views of Nature in Art due: April 13, 2018

This assignment is based on your analysis of and reactions to the Martha Redbone performance we will view on April 3, together with the Sherman Alexie story “One Good Man” that we will discuss in class on Wednesday, April 4. (Note: If you are unable to attend the performance, you can focus your essay on the story alone.)

Assignment

You should compose an essay addressing the following questions.

1. One theme of both Martha Redbone’s performance and of the Alexie story is the desire to create or recover an authentic personal identity based on Native American ancestry. In what ways was that desire expressed in Martha Redbone’s performance and/or in Alexie’s short story. Discuss in detail, citing aspects of the performance and/or the story that are devoted to this theme. Do you think that either attempt to find that identity was entirely successful? What criteria would you use to make a judgment about whether it was successful?

2. On what do you base your own sense of personal identity? Is it the history of your family? Is it the activities you do or the fact that you are now a student at Holy Cross? Is it what you are studying at the college? If it is your family history, how much do you know about that and how did you learn about it? If it is not that family history, what else do you use to create your sense of who you are?

3. Can a person create an individual personal identity independent of where they and their family come from? Can or should we attempt to escape from the past and strike out on our own?

Most of these questions are asking at least partly for personal opinions. So while you should try as much as possible to justify what you are saying by citing details from the story and/or the performance as evidence for the first one, expressing yourself using the first person and a moderately informal style is perfectly OK.

Important Note: It is also perfectly OK with me if you want to share a rough draft of your essay with other students in the class, or with me, to get comments and suggestions. “Testing out” your ideas that way is often an important part of getting them down on paper in an effective form.

Prepare your essay using Microsoft Word or similar software and submit as a .doc file attachment to this email address:

                           jlittle@holycross.edu

Your assignment should arrive in my inbox no later than 5:00pm on Friday, April 13. Aim for around 3 double-spaced pages. (Going a bit short or long is OK, though, as long as you have addressed the questions above adequately and succinctly.)

For this assignment, your are not expected or required to refer to any sources other than “One Good Man.” But if you do, please include a complete listing of all sources you consulted in a numbered References section at the end. If you take a direct quotation from one of them, identify it with an in-line reference of the form “(number, page)” where the number refers to the list of references at the end. Include a page number for printed sources; if you use online sources, that part can be omitted since there is usually not anything corresponding to a physical page number anyway(!)